OpenStudy (anonymous):

as t approaches zero (1/t * sqrt 1+t)-(1/t)

OpenStudy (zzr0ck3r):

\[\frac{ 1 }{ t \sqrt{1+t} }-\frac{ 1 }{ t }\]?

OpenStudy (anonymous):

yes!

OpenStudy (zzr0ck3r):

multiply the 2nd(top and bottom) term by sqrt(1+t) we have (1-sqrt(1-t))/(t*sqrt(t-1)) ok?

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

then devide every term t ((1/t)-(sqrt(1+t)/t)) /((t/t)sqrt(t+1)) = ((1/t)-sqrt(1+t)/t)/sqrt(t+1) ok?

OpenStudy (zzr0ck3r):

ps those sqrt(t-1) should be sqrt(t+1)

OpenStudy (anonymous):

ye i was just about to ask why you used (t-1)

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (zzr0ck3r):

you will stall have 2t on the bottom

OpenStudy (zzr0ck3r):

take it twice:)

OpenStudy (phi):

\[\frac{ 1 }{ t \sqrt{1+t} }-\frac{1}{ t }\] add the fractions to get \[ \frac{1-\sqrt{1+t}}{t\sqrt{1+t}} \] as t ->0 this fraction approaches 0/0 so we can use L'Hopitals Rule First, clear the sort by multiplying top and bottom by sqrt(1+t) \[ \frac{\sqrt{1+t}-1-t}{t^2+t} \] now take the derivative of top and bottom and let t-->0

OpenStudy (zzr0ck3r):

have you learned l`hopital's Rule?

OpenStudy (zzr0ck3r):

@steel11 ?

OpenStudy (anonymous):

nope

OpenStudy (zzr0ck3r):

hmm

OpenStudy (phi):

the bottom give 2t+1 --> 1 as t-->0 the top gives (1/2)(1+t)^(-1/2) -1 -->1/2-1 = -1/2

OpenStudy (zzr0ck3r):

yeah, i made a mistake, but he has not learned L'Hapital's law yet.

OpenStudy (zzr0ck3r):

rule*

OpenStudy (anonymous):

thanks both of you for the help. the answer is -.5.

OpenStudy (phi):

I would be curious to see how to do it without l'hopital.

OpenStudy (zzr0ck3r):

yeah no idea

OpenStudy (zzr0ck3r):

(1-sqrt(x+1))/(x*sqrt(x+1)) top and bottom by (1+sqrt(x+1)) 1-(x+1)/(x*sqrt(x+1)+x(x+1)) devide every term by x -1/(sqrt(x+1)+x+1) run limit =-1/2 @phi @steel11

OpenStudy (zzr0ck3r):

3rd line should be (1-(x+1)/(x*sqrt(x+1)+x(x+1))

OpenStudy (phi):

looks good